Desireemaya:

Isaac applied the steps below to find the product of (4.2)(–5.4).Step 1: (4.2)(–5.4) = (–5.4)(4.2) Step 2: = (–5.4)(4) + (–5.4)(0.2) Step 3: = (–21.6) + (–1.08) Step 4: = –22.68

jhonyy9:

@XioGonz what you think about this ?

jhonyy9:

ty

XioGonz:

(4.2) (-54) = (-54) (4.2) = (-5.4) (4) + (-5.4) (0.2) ((-5.4) (4.2) = (-5.4) (4 × 0.2) = (-5.4) (4) + (-5.4) (0.2)) = (-21.6) + (-1.08) = 22.68 We observe that in Step 2: (-5.4) (4.2) = (-5.4) (4 × 0.2) = (-5.4) (4) + (-5.4) (0.2) So which part did he apply the distributive property to?